OpenStudy (anonymous):

using the point slope form write an equation for the line with slope -2 containing point (3, 4) please help!

OpenStudy (chaise):

y = mx + b Your y is 4 your x is 3 your m is -2 You can solve for b now 4 = -2x3 + b 4 = -6 + b b = 10 y = mx +b y = -2x+10